In a bag of jellybeans, 1/3 are cherry and 1/4 are licorices. If the remaining 20 jellybeans are orange, how many jellybeans are in the bag?
Answer by rfer(16322)

1/3+1/4=
4/12+3/12=7/12
20 jellybeans=5/12
20/(5/12)=
20*12/5=48 total